APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday March 6, 2007 (MJD 54166) ---=== OBSERVING TEAM ===--- Howard Brewington Dan Oravetz Suvrath Mahadevan (UF) Scott Fleming (UF) Audrey Simmons ---=== OBSERVING PLAN ===--- ET Science! ---=== OBSERVING SUMMARY ===--- The IOP sessions were hanging this afternoon, so Jon Brinkmann rebooted sdsshost2 and things proceeded to work normally thereafter. We observed 8 different ET plates and took 11 exposures. We started the DIMM up and have been using it for the past four nights and haven't had any problems. We also used the test version of IOP the last four nights and haven't encountered any problems. Altitude slip and Rotator aborts: When slewing from Cartridge 8 to the instrument change position (09:58Z), we had an altitude slip (alt~70). During the same slew we had a rotator abort which occurred when the rotator was at rota=251. There was no indication of the ET fiber contacting the telescope to cause the abort. NOTE: This is the third consecutive night where we had a rotator abort when slewing from Cartridge 8 to the instrument change position (and all within about 2 minutes of the same time). Last night and tonight the rotator was at rota=251 when the abort occurred (we don't have a record of the rotator position from two nights ago, but we think it was ~251 degrees as well). At 11:02Z, the rotator aborted several times when slewing from the instrument change position to the Cartridge 1 field. The rotator never even began to move while the telescope was slewing in alt and az to the field. After several attempts, the rotator responded. Again, there was no indication that the ET fiber was contacting the telescope to cause the abort. ---=== OBSERVING LOG ===--- Afternoon: --------- The IOP sessions were hanging, so Jon Brinkmann rebooted sdsshost2. Afternoon checkout went smoothly. Night: ----- 01:40Z - Started with Cartridge 6, plate E8016. Needed an fk5. One 50-minute exposure with the iodine cell in place (data). 02:51Z - Cartridge 2, plate E8018. GSOGTF One 50-minute exposure with the iodine cell in place (data). 03:53Z - Cartridge 9, plate E8017. Needed an fk5 and cleared the offsets to find guide stars. One 50-minute exposure with the iodine cell in place (data). One 50-minute exposure with the iodine cell out (template). 05:53Z - Cartridge 3, plate E8024. GSOGTF One 50-minute exposure with the iodine cell out (template). One 50-minute exposure with the iodine cell in place (data). 07:48Z - Cartridge 5, plate E8025. GSOGTF One 50-minute exposure with the iodine cell in place (data). 08:56Z - Cartridge 8, plate E8021. Needed an fk5. One 50-minute exposure with the iodine cell in place (data). 09:58Z - When slewing from Cartridge 8 to the instrument change position we had an altitude slip (at ~alt=70). 09:58Z - When slewing from Cartridge 8 to to the instrument change position we also had a rotator abort. The abort occurred when the rotator was at rota=251. There was no indication of the ET fiber contacting the telescope to cause the abort. NOTE: For the third consecutive night, the rotator has aborted when slewing from Cartridge 8 to the instrument change position (all within 2 minutes of the same Zulu time as well). Last night and tonight, the rotator was at 251 degrees when it aborted (we don't have a record of the rotator position from two nights ago, but we think it was also 251 degrees). 10:10Z - Cartridge 7, plate E8022. GSOGTF One 50-minute exposure with the iodine cell in place (data). 11:02Z - Rotator aborted several times when slewing from the instrument change position to the Cartridge 1 field. The rotator never even began to move while the telescope was slewing in alt and az to the field. After several attempts, the rotator responded. Again, there was no indication that the ET fiber was contacting the telescope to cause the abort. 11:12Z - Cartridge 1, plate E8026. Needed an fk5. One 50-minute exposure with the iodine cell in place (data). One 30-minute exposure with the iodine cell in place (data). ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): N/A QA Procedures Done (y/n): N/A UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 43236, plate -9999 ------- 23:47 43236 0.0 bias 23:50 43237 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 01:50Z 6 8016 26.05 -0.0003 54.15 0.0030 201.64 -0.0076 1.000160 02:53Z 2 8018 174.84 0.0149 81.80 0.0017 -5.74 0.0298 1.000160 03:55Z 9 8017 -11.75 -0.0003 62.00 0.0002 170.09 0.0027 1.000160 05:54Z 3 8024 19.02 -0.0016 74.67 0.0016 196.74 0.0314 1.000120 07:54Z 5 8025 -1.80 -0.0016 60.03 -0.0004 178.43 -0.0148 1.000080 09:01Z 8 8021 95.06 0.0005 50.64 0.0036 251.04 -0.0120 1.000180 10:23Z 7 8022 111.23 0.0006 62.82 0.0034 -88.37 -0.0132 1.000180 11:18Z 1 8026 212.90 0.0022 62.61 0.0003 50.32 -0.0132 1.000110 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:50Z 6 8016 1.00016 -1580 -65 75 26 54.1 7.4 7 19 ---- 1.7 02:53Z 2 8018 1.00016 -1580 -166 25 175 81.8 6.3 3 64 ---- 1.4 03:55Z 9 8017 1.00016 -1580 -161 0 -12 62.0 6.3 2 120 ---- 1.3 05:54Z 3 8024 1.00012 -1185 118 -25 19 74.7 6.7 2 274 ---- 1.0 07:54Z 5 8025 1.00008 -790 523 25 -2 60.0 5.9 1 162 ---- 1.4 09:01Z 8 8021 1.00018 -1778 -292 0 95 50.6 6.0 6 18 ---- 1.5 10:23Z 7 8022 1.00018 -1778 -300 25 111 62.8 5.4 6 20 ---- 1.5 11:17Z 1 8026 1.00011 -1086 247 0 213 62.7 6.6 11 13 ---- 1.3 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:25Z 49 -5 3 268 (W) 378 - 23:56Z 49 -5 3 129 (SE) 384 - 00:26Z 48 -5 3 44 (NE) 385 - 00:56Z 46 -4 6 24 (NNE) 399 - 01:29Z 45 -2 8 11 (N) 449 - 02:00Z 45 -1 6 19 (NNE) 507 - 02:30Z 43 0 4 21 (NNE) 540 - 03:01Z 44 0 1 130 (SE) 496 - 03:33Z 44 1 1 173 (S) 478 - 04:07Z 44 0 2 170 (S) 432 - 04:39Z 45 -1 2 255 (WSW) 454 - 05:11Z 45 5 4 271 (W) 697 - 05:43Z 45 6 2 248 (WSW) 708 - 06:15Z 44 7 1 254 (WSW) 698 - 06:48Z 43 7 0 252 (WSW) 684 - 07:20Z 43 5 1 178 (S) 698 - 07:50Z 43 8 1 73 (ENE) 669 - 08:23Z 43 9 3 32 (NNE) 791 - 08:54Z 43 8 2 48 (NE) 717 - 09:24Z 42 7 3 35 (NE) 780 - 09:55Z 43 8 3 18 (NNE) 779 - 10:27Z 41 8 7 14 (NNE) 836 - 10:59Z 41 8 10 8 (N) 858 - 11:32Z 43 5 11 9 (N) 903 - 12:05Z 42 5 6 9 (N) 774 - 12:37Z 42 10 6 14 (NNE) 575 - ---=== TELESCOPE STATUS ===--- 00:34Z - Began ventilation. Doors and louvers open, fans on. 01:05Z - Enclosure off. 12:40Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 12:55Z Telescope stowed at: 30,121,00 mount Instrument mounted: ET Cartridge 1 Counterweights at: 245 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 173 lbs, 20 psi Imager 144 lbs, 18 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_31_0 Watcher: v2_35_0 MCP: v6_3_2 TPM: tpm_v3_5_0 AstroDa: v15_13 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0010 -0.0050 Axial B 0.0030 0.0090 Axial C 0.0150 0.0160 Trans D 0.0030 0.0040 Lateral E 9.9695 9.9568 Lateral F 11.5570 11.5443 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 9.7 Alt.: 30.258484 MIGS TONIGHT NOMINAL Axial A -0.0280 -0.0240 Axial B -0.0320 -0.0280 Axial C -0.0410 -0.0370 Trans D -0.0120 -0.0040 GALILS Commanded: 1614043. 1594704. 1543101. -10550. -5800. Actual: 1614050. 1594700. 1543100. -10550. -5800. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1257.00 19.99 -10.00 -100.43 176.57 ---=== PROBLEMS IN DETAIL ===--- Miscellaneous Watcher report: altitude (Failed to read MEI data for altitude: Move distance is zero (2007-03-06 23:59:33Z sdssmcp): 2007-03-06 23:59:46Z ) azimuth (Failed to read MEI data for azimuth: Move distance is zero (2007-03-06 23:59:33Z sdssmcp): 2007-03-06 23:59:46Z ) rotator (Failed to read MEI data for rotator: Move distance is zero (2007-03-06 23:59:33Z sdssmcp): 2007-03-06 23:59:46Z ) dewar (Imager Dewar Low Weight:: 2007-03-07 03:29:06Z) env temp (Excessive Return - Ambient Difference:: 2007-03-07 01:10:44Z) tracking (Excessive Az Tracking Error:: 2007-03-07 01:09:39Z) currents (High Rot Drive Current:: 2007-03-07 01:36:30Z) start move failed; offsetting and trying again (2007-03-07 00:00:28Z sdssmcp) (start_move failed; offsetting and trying again (2007-03-07 00:00:28Z sdssmcp): 2007-03-07 00:00:40Z ) failed to take semMEI: S objLib OBJ TIMEOUT (3997700) (2007-03-07 09:50:02Z sdssmcp) (failed to take semMEI: S_objLib_OBJ_TIMEOUT (3997700) (2007-03-07 09:50:02Z sdssmcp): 2007-03-07 09:50:12Z ) Windscreen touch 09:57Z.